# An alternating low-rank projection approach for partial differential equations with random inputs

### v1

@@ 4 Numerical results
In this section, we consider two problems: a diffusion problem and a Helmholtz problem. For the diffusion problem, we employ the standard finite element method to generate the matrices $\bm{A}_{i}$, where $i\in\{0,1,\ldots,K\}$. For the Helmholtz problem, we utilize the codes associated with [34] for convenience. All the results presented in this study are obtained using MATLAB R2016a on a desktop computer with 2.90GHz Intel Core i7-10700 CPU.
@@ 4.1 Test problem 1
$-\nabla\cdot(a(\bm{x},\bm{\xi})\nabla u(\bm{x},\bm{\xi}))=1&\quad\mbox{in}\quad D\times\Gamma,\\ u(\bm{x},\bm{\xi})=0&\quad\mbox{on}\quad\partial D\times\Gamma,\\ \end{aligned}\right.$
@@
where $\{\lambda_{i},c_{i}(\bm{x})\}_{i=1}^{N}$ are the eigenpairs of $\mathrm{Cov}\,(\bm{x},\bm{y})$, $\{\xi_{i}\}^{N}_{i=1}$ are uncorrelated random variables, and $N$ is the number of KL modes retained. For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are identically independent distributed uniform random variables on $[-1,1]$.
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (19) and (23)-(24), are solved using the preconditioned conjugate gradients (PCG) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
@@
In Figure 2, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (25). The ‘optimal’ solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (18) for it deteriorates for smaller values of $\mathtt{RelTol}$.
@@ 4.2 Test problem 2
For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are independent and uniformly distributed within the range $[-1,1]$. We use the perfectly matched layers (PML) to simulate the Sommerfeld condition [36], and generate the matrices $\{\bm{A}_{i}\}_{i=1}^{K}$ using the codes associated with [34].
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (19) and (23)-(24), are solved using the preconditioned bi-conjugate gradient stabilized (Bi-CGSTAB) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
@@
In Figure 5, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (25). The ‘optimal’ solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (18) for it deteriorates for smaller values of $\mathtt{RelTol}$.

### v2

@@ 4 Numerical results
In this section, we consider two problems: a diffusion problem and a Helmholtz problem. For both test problems, spatial discretization is performed using $Q_{1}$ finite elements based on the IFISS package [39]. All the results presented in this study are obtained using MATLAB R2016a on a desktop computer with 2.90GHz Intel Core i7-10700 CPU.
@@ 4.1 Test problem 1
[…]
@@
where $\{\lambda_{i},c_{i}(\bm{x})\}_{i=1}^{N}$ are the eigenpairs of $\mathrm{Cov}\,(\bm{x},\bm{y})$, $\{\xi_{i}\}^{N}_{i=1}$ are uncorrelated random variables, and $N$ is the number of KL modes retained. For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are identically independent distributed uniform random variables on [1,1][-1,1].
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (24) and (30)–(32), are solved using the preconditioned conjugate gradients (PCG) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
@@
In Figure 2, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (34). The ‘optimal’ low-rank approximation of the solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (20) for it deteriorates for smaller values of $\mathtt{RelTol}$.
@@
To further evaluate the efficiency of the proposed methods, we compare it with the low-rank solver developed in [16], which employs an SVD-based truncation operator $\mathcal{T}_{\mathtt{RelTol}}$, referred to as LRSSVD in the following discussion. In this test problem, the following truncation strategy is adopted: truncation is applied before preconditioning, after preconditioning, after the computation of the matrix-vector product (performed in low-rank form), and after every summation during the iterative process of PCG.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 4: Relative errors with respect to rank $k$.
Figure 4 shows the relative errors with respect to the rank $k$. For the AltLRP approach, the relative error is almost indistinguishable from that of the ‘optimal’ low-rank approximation of the solution. In comparison, the LRSSVD approach requires slightly higher ranks to achieve the same level of accuracy.
The SimLRP approach, however, exhibits a faster growth in rank compared to both the AltLRP and LRSSVD methods—particularly when the stochastic dimension is $N=7$ or $N=10$. In these high-dimensional cases, the SimLRP method shows limited improvement in accuracy as the rank increases, once the relative error reaches approximately $10^{-5}$.
Another noteworthy observation is that, across all methods, the curves of relative error versus rank $k$ remain nearly unchanged for different mesh sizes in the physical space. Note that the relative error of $\bm{U}$, defined in (34), can be regarded as a discretized version of the relative error of $u^{p}(\bm{x},\bm{\xi})$ defined in (17). This behavior is, in fact, a direct consequence of Theorem 3.2.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 5: Relative errors with respect to CPU time in seconds for the LRSSVD approach.
Figure 5 shows the CPU time in seconds for the LRSSVD approach. From Figure 3 and 5, it is clear that the AltLRP approach requires significantly less CPU time than the LRSSVD approach to achieve a given accuracy. In addition, if the required relative error is above $10^{-5}$, the SimLRP approach can also be applied.
The efficiency of the AltLRP approach stems from two main reasons. First, the method proposed in this work determines the rank required for a given accuracy by analyzing the singular values of quasimatrices associated with the truncated gPC expansion of the solution on a coarse physical grid, allowing the construction of a fixed-rank approximation. This avoids the need for truncation during the PCG iterations for solving the linear systems. Second, while the LRSSVD method operates directly on the full linear system and applies low-rank techniques during the iteration, the proposed AltLRP approach constructs a much smaller projected linear system based on the identified low-rank subspace. This projection significantly reduces computational cost compared to the LRSSVD approach.
It should be pointed out that the computational time in LRSSVD depends on several factors, such as the method used for SVD updates and the termination criteria of the iteration. In this work, we only employ a simple implementation of  [16], which may not achieve its optimal efficiency.
@@ 4.2 Test problem 2
For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are independent and uniformly distributed within the range [1,1][-1,1]. We use the perfectly matched layers (PML) to simulate the Sommerfeld condition [40], and discretize the physical space using $Q_{1}$ finite elements based on the IFISS package [39].
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (24) and (30)–(32), are solved using the preconditioned bi-conjugate gradient stabilized (Bi-CGSTAB) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
@@
In Figure 7, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (34). The ‘optimal’ low-rank approximation of the solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (20) for it deteriorates for smaller values of $\mathtt{RelTol}$.
@@
To further evaluate the efficiency of the proposed method, we compare it with the LRSSVD approach developed in [16]. The truncation strategy used here is the same as in the previous test problem: truncation is applied before preconditioning, after preconditioning, after the matrix-vector product (performed in low-rank form), and after each summation during the iterative process of Bi-CGSTAB.
It is worth noting that although this test problem involves only $N$ random variables, the number of terms on the left-hand side of equation (10) is $K=(N+1)^{2}+1$, which is significantly larger than that in Test Problem 1. This leads to increased memory requirements during the iterative process, since the rank after the matrix-vector product (in low-rank form) becomes $K\times k$. As a result, some cases run out of memory.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 9: Relative errors with respect to rank $k$.
Figure 9 shows the relative errors with respect to the rank $k$. For the LRSSVD approach, the relative error is almost indistinguishable from that of the ‘optimal’ low-rank approximation of the solution. In comparison, the AltLRP approach requires slightly higher ranks to achieve the same level of accuracy. The SimLRP approach, however, exhibits a faster growth in rank compared to both the AltLRP and LRSSVD methods—particularly when the stochastic dimension is $N=7$ or $N=10$. In these high-dimensional cases, the SimLRP method shows limited improvement in accuracy as the rank increases, once the relative error reaches approximately $10^{-5}$.
Another noteworthy observation is that, across all methods, the curves of relative error versus rank $k$ remain nearly unchanged for different mesh sizes in the physical space. Note that the relative error of $\bm{U}$, defined in (34), can be regarded as a discretized version of the relative error of $u^{p}(\bm{x},\bm{\xi})$ defined in (17). This behavior is, in fact, a direct consequence of Theorem 3.2.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 10: Relative errors with respect to CPU time in seconds for the LRSSVD approach.
Figure 10 shows the CPU time in seconds for the LRSSVD approach. Note that for the case of $N=10$ and $n_{x}=257^{2}$, the computation runs out of memory when the tolerance $\mathtt{RelTol}$ is set to $10^{-5}$, $10^{-6}$, and $10^{-7}$. From Figure 8 and 10, it is clear that the AltLRP approach requires significantly less CPU time than the LRSSVD approach to achieve a given accuracy. In addition, if the required relative error is above $10^{-5}$, the SimLRP approach can also be applied.